Consider the combustion reaction for octane
(
C
8
H
18
)
,
which is a primary component of gasoline.

2
C
8
H
18
+
25
O
2

16
CO
2
+
18
H
2
O

How many moles of
CO
2
are emitted into the atmosphere when
12.1
g
C
8
H
18
is burned?
1
Expert's answer
2019-11-26T06:39:48-0500

.Solution.

n(C8H18) = 12.1/114 = 0.106 moles

0.106 - x

2 - 16

n(CO2) = 0.106*16/2= 0.849 moles

Answer. 0.849 moles